# Bounding the SNPR distance between two tree-child networks using generalised agreement forests

### v1

@@ 3. Phylogenetic digraphs and their extensions
In this section, we provide formal definitions of the concepts of phylogenetic digraphs and their extensions. As we will see in Section 5, these definitions generalise agreement forests for two phylogenetic trees to two phylogenetic networks. We start by providing some high-level ideas for the upcoming definitions that may guide the reader in developing some intuition. Essentially, a phylogenetic digraph ${\mathcal{D}}$ of a phylogenetic network ${\mathcal{N}}$ on $X$ is a collection of acyclic digraphs whose vertices of out-degree zero are bijectively labelled with the elements in $X$ and for which there is a vertex-disjoint embedding of its components in ${\mathcal{N}}$. Let ${\mathcal{M}}$ be such an embedding of ${\mathcal{D}}$ in ${\mathcal{N}}$, and let $v$ be a vertex of ${\mathcal{M}}$ that either has in-degree zero, or is a reticulation of ${\mathcal{N}}$ and has in-degree one and out-degree one. We obtain an extension ${\mathcal{R}}$ of ${\mathcal{M}}$ by starting at $v$ and extending ${\mathcal{M}}$ towards the root by adding edges of ${\mathcal{N}}$ in a certain algorithmic way and then repeating this process for all such vertices of ${\mathcal{M}}$. As suggested by the phrase extension of ${\mathcal{M}}$, ${\mathcal{R}}$ contains all edges of ${\mathcal{M}}$. Although ${\mathcal{D}}$ may have several embeddings in ${\mathcal{N}}$, each embedding is anchored at the leaves of ${\mathcal{N}}$ due to the requirement that the leaves of ${\mathcal{D}}$ bijectively map to the elements in $X$. As we will see in Section 4, for the purpose of computing the minimum number of edges in ${\mathcal{N}}$ that are not contained in an extension relative to a given phylogenetic digraph ${\mathcal{D}}$ of ${\mathcal{N}}$, where the minimum is taken over all embeddings of ${\mathcal{D}}$ in ${\mathcal{N}}$ and their extensions, it is sufficient to consider only a single extension of ${\mathcal{D}}$.
Phylogenetic digraphs. Let $D$ be a connected directed graph, and let $Y$ be a non-empty finite set. We say that $D$ is a leaf-labelled acyclic digraph on $Y$ if it has at most one vertex $\rho$ of in-degree zero and out-degree one, its leaves with in-degree one and out-degree zero are bijectively labelled with the elements in $Y$, and all other vertices have in-degree zero and out-degree two, in-degree one and out-degree two, or in-degree two and out-degree one. If $|Y|=1$, then the isolated vertex labelled with the element in $Y$ is also considered to be a leaf-labelled acyclic digraph on $Y$. Similar to the leaf set of a phylogenetic network, we refer to $Y$ as the leaf set of $D$ and denote it by ${\mathcal{L}}(D)$. In contrast to a phylogenetic network, a leaf-labelled acyclic digraph $D$ may have more than one vertex with in-degree zero. Moreover, for a vertex $w$ in $D$ that has two parents $v$ and $v^{\prime}$, there does not necessarily exist a vertex $u$ such that there are edge-disjoint directed paths from $u$ to $v$ and from $u$ to $v^{\prime}$ in $D$.
Let ${\mathcal{N}}$ be a phylogenetic network on $X$ with root $\rho$, and let $D$ be a leaf-labelled acyclic digraph on $Y$ with $Y\subseteq X\cup\{\rho\}$. We say that ${\mathcal{N}}$ displays $D$ if there exists a subgraph of ${\mathcal{N}}$ that is isomorphic to $D$ up to suppressing vertices with in-degree one and out-degree one, in which case we call the subgraph an embedding $M$ of $D$ in ${\mathcal{N}}$ and view the edge set of $M$ as a subset of the edge set of ${\mathcal{N}}$. More generally, for a collection ${\mathcal{D}}=\{D_{1},D_{2},\ldots,D_{k}\}$ of leaf-labelled acyclic digraphs, we say that ${\mathcal{N}}$ displays ${\mathcal{D}}$ if there exists an embedding $M_{i}$ of $D_{i}$ in ${\mathcal{N}}$ for each $i\in\{1,2,\ldots,k\}$ such that $M_{j}$ and $M_{j^{\prime}}$ are vertex disjoint for all distinct $j,j^{\prime}\in\{1,2,\ldots,k\}$, in which case we refer to =${\mathcal{M}}=\{M_{1},M_{2},\ldots,M_{k}\}$ as an embedding of ${\mathcal{D}}$ in ${\mathcal{N}}$. Now let ${\mathcal{M}}$ be an embedding of ${\mathcal{D}}$ in ${\mathcal{N}}$. Recalling that we allow tree vertices in ${\mathcal{M}}$ to have in-degree and out-degree one, we say that ${\mathcal{M}}$ is tree-child if each non-leaf vertex of ${\mathcal{M}}$ has a child that is a tree vertex or a leaf.
Let ${\mathcal{N}}$ be a phylogenetic network on $X$. Let ${\mathcal{D}}=\{D_{\rho},D_{1},D_{2},\ldots,D_{k}\}$ be a collection of leaf-labelled acyclic digraphs. Then ${\mathcal{D}}$ is called a phylogenetic digraph of ${\mathcal{N}}$ if the following three properties are satisfied:
@@
    the leaf sets ${\mathcal{L}}(D_{\rho}),{\mathcal{L}}(D_{1}),{\mathcal{L}}(D_{2}),\ldots,{\mathcal{L}}(D_{k})$ partition $X\cup\{\rho\}$ and $\rho\in{\mathcal{L}}(D_{\rho})$,
@@
    there exists an embedding =${\mathcal{M}}=\{M_{\rho},M_{1},M_{2},\ldots,M_{k}\}$ of ${\mathcal{D}}$ in ${\mathcal{N}}$.
@@
[…]
Figure 1. Two phylogenetic networks ${\mathcal{N}}$ and ${\mathcal{N}}^{\prime}$ and three phylogenetic digraphs ${\mathcal{D}}_{1}$, ${\mathcal{D}}_{2}$, and ${\mathcal{D}}_{3}$ for ${\mathcal{N}}$ and ${\mathcal{N}}^{\prime}$. For each $i\in\{1,2,3\}$, ${\mathcal{R}}_{i}$ is an extension of ${\mathcal{D}}_{i}$ in ${\mathcal{N}}$ and ${\mathcal{R}}_{i}^{\prime}$ is an extension of ${\mathcal{D}}_{i}$ in ${\mathcal{N}}^{\prime}$, where the edges of the extensions are indicated by thick lines in ${\mathcal{N}}$ and ${\mathcal{N}}^{\prime}$, respectively. Note that ${\mathcal{R}}_{3}$ is not a root extension of ${\mathcal{D}}_{3}$ in ${\mathcal{N}}$.
Extensions and root extensions. Let ${\mathcal{N}}$ be a phylogenetic network on $X$. Furthermore, let =${\mathcal{M}}=\{M_{\rho},M_{1},M_{2},\ldots,M_{k}\}$ be an embedding of a phylogenetic digraph ${\mathcal{D}}=\{D_{\rho},D_{1},D_{2},\ldots,D_{k}\}$ of ${\mathcal{N}}$. We obtain an extension ${\mathcal{R}}$ of ${\mathcal{D}}$ in ${\mathcal{N}}$ from ${\mathcal{M}}$ by initially setting ${\mathcal{R}}={\mathcal{M}}$ and then repeatedly applying one of the following two operations until no further such operation is possible:
@@
Observe that ${\mathcal{R}}$ contains exactly $k+1$ connected components and that there is a natural bijection between these components and the components in ${\mathcal{D}}$. We therefore set =${\mathcal{R}}=\{R_{\rho},R_{1},R_{2},\ldots,R_{k}\}$ and call $R_{i}$ an extension of $D_{i}$ in ${\mathcal{N}}$ for each $i\in\{\rho,1,2,\ldots,k\}$. It follows from the construction of ${\mathcal{R}}$ that there is no vertex of out-degree two in ${\mathcal{R}}$ that is not also a vertex of out-degree two in ${\mathcal{M}}$. Moreover, by construction, any underlying cycle in ${\mathcal{R}}$ is also an underlying cycle in ${\mathcal{M}}$. Lastly, we define ${\mathcal{R}}$ to be tree-child precisely if ${\mathcal{M}}$ is tree-child, and refer to ${\mathcal{M}}$ as the embedding of ${\mathcal{D}}$ that underlies ${\mathcal{R}}$.
@@
We next introduce a special type of extension. We call an extension ${\mathcal{R}}$ of ${\mathcal{D}}$ in ${\mathcal{N}}$ a root extension of ${\mathcal{D}}$ in ${\mathcal{N}}$ if it can be obtained from ${\mathcal{M}}$ by initially setting ${\mathcal{R}}={\mathcal{M}}$ and then repeatedly applying (E1) only until no further such operation is possible. Now let =${\mathcal{R}}=\{R_{\rho},R_{1},R_{2},\ldots,R_{k}\}$ be a root extension of ${\mathcal{D}}$ in ${\mathcal{N}}$. Similar to the terminology for an extension, we call $R_{i}$ a root extension of $D_{i}$ in ${\mathcal{N}}$ for each $i\in\{\rho,1,2,\ldots,k\}$. Let $r$ be a vertex of in-degree zero and out-degree zero or two in ${\mathcal{D}}$, and let $P$ be the unique maximal length directed path in ${\mathcal{R}}$ that starts at a vertex $u$ of in-degree zero and ends at $r$. We refer to $P$ as the root path of $r$. Note that $P$ may have no edge in which case $u=r$. If $u\neq r$, then $u$ has in-degree zero and out-degree one in ${\mathcal{R}}$. Figure 1 illustrates the concepts of phylogenetic digraphs, extensions, and root extensions.
@@
To ease reading throughout the remainder of the paper, we often consider a phylogenetic network ${\mathcal{N}}$, a phylogenetic digraph ${\mathcal{D}}$ of ${\mathcal{N}}$, and an extension ${\mathcal{R}}$ of ${\mathcal{D}}$ in ${\mathcal{N}}$. In this case, we view the vertex and edge set of ${\mathcal{R}}$ (as well as the vertex and edge set of any embedding of ${\mathcal{D}}$ in ${\mathcal{N}}$) as a subset of the vertex and edge set of ${\mathcal{N}}$, respectively. Furthermore, for clarity, the in-degree (resp. out-degree) of a vertex $v$ in ${\mathcal{R}}$ refers to the number of edges in ${\mathcal{R}}$ that are directed into (resp. out of) $v$. Lastly, let ${\mathcal{D}}=\{D_{\rho},D_{1},D_{2},\ldots,D_{k}\}$ be a phylogenetic digraph of a phylogenetic network ${\mathcal{N}}$. Then there exists an embedding =${\mathcal{M}}=\{M_{\rho},M_{1},M_{2},\ldots,M_{k}\}$ of ${\mathcal{D}}$ in ${\mathcal{N}}$ such that $M_{i}$ and $M_{j}$ are vertex disjoint for all distinct $i,j\in\{\rho,1,2,\ldots,k\}$ and an extension =${\mathcal{R}}=\{R_{\rho},R_{1},R_{2},\ldots,R_{k}\}$ of ${\mathcal{D}}$ in ${\mathcal{N}}$ such that $R_{i}$ and $R_{j}$ are vertex disjoint for all distinct $i,j\in\{\rho,1,2,\ldots,k\}$. It follows that each edge in ${\mathcal{D}}$ corresponds to a unique directed path in ${\mathcal{M}}$ (resp. ${\mathcal{R}}$) whose non-terminal vertices all have in-degree one and out-degree one in ${\mathcal{M}}$ (resp. ${\mathcal{R}}$), and each vertex in ${\mathcal{D}}$ corresponds to a unique vertex in ${\mathcal{M}}$ (resp. ${\mathcal{R}}$). Reversely, each edge in ${\mathcal{M}}$ corresponds to a unique edge in ${\mathcal{D}}$. We will freely use this correspondence throughout the paper.
4. Properties of extensions
@@
Let ${\mathcal{D}}=\{D_{\rho},D_{1},D_{2},\ldots,D_{k}\}$, and let =${\mathcal{R}}=\{R_{\rho},R_{1},R_{2},\ldots,R_{k}\}$. To see that (i) holds, recall that ${\mathcal{R}}$ does not contain any vertex with out-degree zero that is not in $X\cup\{\rho\}$. To complete the proof, we establish (ii). By definition of ${\mathcal{D}}$, the root and each leaf of ${\mathcal{N}}$ is contained in ${\mathcal{R}}$. Towards a contradiction, we may therefore assume that there is a tree vertex or a reticulation in ${\mathcal{N}}$ that is not contained in ${\mathcal{R}}$. Let $v$ be a vertex of ${\mathcal{N}}$ that is not in ${\mathcal{R}}$ such that every vertex that is distinct from $v$ and lies on a directed path from $v$ to a leaf in ${\mathcal{N}}$ is in ${\mathcal{R}}$. Furthermore, let $w$ be a child of $v$. If $w$ is a tree vertex or leaf in ${\mathcal{N}}$, then there exists a component $R_{i}$ with $i\in\{\rho,1,2,\ldots,k\}$ such that $w$ has in-degree zero in $R_{i}$, thereby contradicting that $R_{i}$ is an extension of $D_{i}$ as we can apply (E1). Otherwise, if $w$ is a reticulation in ${\mathcal{N}}$, then it follows from (i) that there exists a component $R_{i}$ with $i\in\{\rho,1,2,\ldots,k\}$ such that $w$ has in-degree zero, or in-degree one and out-degree one in $R_{i}$, thereby again contradicting that $R_{i}$ is an extension of $D_{i}$ as we can apply either (E1) or (E2), respectively. ∎
Lemma 4.2.
@@
Let $u$ be the parent of $v$, and let $w$ and $w^{\prime}$ be the two children of $v$ in ${\mathcal{N}}$. Furthermore, let =${\mathcal{R}}=\{R_{\rho},R_{1},R_{2},\ldots,R_{k}\}$. We first show that (i) holds. Since ${\mathcal{M}}$ does not contain any vertex of out-degree zero that is not in $X\cup\{\rho\}$, it follows that at least one of $(v,w)$ and $(v,w^{\prime})$ is in ${\mathcal{M}}$. Without loss of generality, we may assume that $(v,w)\in{\mathcal{M}}$. If $(v,w^{\prime})\in{\mathcal{M}}$, then the result clearly holds. On the other hand, if $(v,w^{\prime})\notin{\mathcal{M}}$, then it follows from the definition of an extension that $(v,w^{\prime})\notin{\mathcal{R}}$, thereby establishing (i). We now turn to (ii). It follows from Lemma 4.1(ii) that $v\in{\mathcal{R}}$. Since ${\mathcal{R}}$ does not contain any vertex with out-degree zero that is not in $X\cup\{\rho\}$, at least one of $(v,w)$ and $(v,w^{\prime})$ is contained in ${\mathcal{R}}$. Moreover, again by the definition of an extension, at most one of $(v,w)$ and $(v,w^{\prime})$ is contained in ${\mathcal{R}}$. The lemma now follows. ∎
@@
Let ${\mathcal{M}}$ and ${\mathcal{M}}^{\prime}$ be the embeddings of ${\mathcal{D}}$ in ${\mathcal{N}}$ that underlie ${\mathcal{R}}$ and ${\mathcal{R}}^{\prime}$, respectively. First assume that ${\mathcal{M}}={\mathcal{M}}^{\prime}$. Since $|E_{\mathcal{N}}-E_{\mathcal{M}}|=|E_{\mathcal{N}}-E_{{\mathcal{M}}^{\prime}}|$ and each iteration of (E1) or (E2) in the construction of ${\mathcal{R}}$ and ${\mathcal{R}}^{\prime}$ either adds a new edge $(v,w)$ such that $w$ is already in ${\mathcal{R}}$ and $v$ is not already in ${\mathcal{R}}$ or a new edge $(v^{\prime},w^{\prime})$ such that $w^{\prime}$ is already in ${\mathcal{R}}^{\prime}$ and $v^{\prime}$ is not already in ${\mathcal{R}}^{\prime}$, the result follows from Lemma 4.1(ii). Second assume that ${\mathcal{M}}\neq{\mathcal{M}}^{\prime}$. Consider an edge $e=(v,w)$ of ${\mathcal{N}}$. If $v=\rho$, then, as $\rho$ has out-degree one in ${\mathcal{N}}$, either $e\in{\mathcal{R}}$ and $e\in{\mathcal{R}}^{\prime}$, or $e\notin{\mathcal{R}}$ and $e\notin{\mathcal{R}}^{\prime}$. Furthermore, if $v$ is a reticulation in ${\mathcal{N}}$, then it follows from Lemma 4.1(i) that $e\in{\mathcal{R}}$ and $e\in{\mathcal{R}}^{\prime}$. We next consider all edges in ${\mathcal{N}}$ that are directed out of a tree vertex. Let $v$ be a tree vertex of ${\mathcal{N}}$, and let $e=(v,w)$ and $e^{\prime}=(v,w^{\prime})$ be the two edges directed out of $v$. We next consider four cases that will subsequently be used for a counting argument.

### v2

@@ 3 Phylogenetic digraphs and their extensions
In this section, we provide formal definitions of the concepts of phylogenetic digraphs and their extensions. As we will see in Section 5, these definitions generalise agreement forests for two phylogenetic trees to two phylogenetic networks. We start by providing some high-level ideas that may guide the reader in developing some intuition before providing formal definitions. Essentially, a phylogenetic digraph $\mathcal{D}$} of a phylogenetic network $\mathcal{N}$} on $X$ and with root $\rho$ is a collection of directed acyclic graphs that each contain at least one element in $X$ with the exception that $\rho$ may be a singleton in $\mathcal{D}$}, whose vertices of out-degree zero are bijectively labelled with the elements in $X$, and for which there is a vertex-disjoint embedding of its components in $\mathcal{N}$}. Let $\mathcal{M}$} be such an embedding of $\mathcal{D}$} in $\mathcal{N}$}, and let $v$ be a vertex of $\mathcal{M}$} that either has in-degree zero, or is a reticulation of $\mathcal{N}$} and has in-degree one and out-degree one. We obtain an extension $\mathcal{R}$} of $\mathcal{M}$} by starting at $v$ and extending $\mathcal{M}$} towards the root by adding edges of $\mathcal{N}$} in a certain algorithmic way and then repeating this process for all such vertices of $\mathcal{M}$}. As suggested by the phrase extension of $\mathcal{M}$}, $\mathcal{R}$} contains all edges of $\mathcal{M}$}. Although $\mathcal{D}$} may have several embeddings in $\mathcal{N}$}, each embedding is anchored at the leaves of $\mathcal{N}$} due to the requirement that the leaves of $\mathcal{D}$} bijectively map to the elements in $X$. As we will see in Section 4, for the purpose of computing the minimum number of edges in $\mathcal{N}$} that are not contained in an extension relative to a given phylogenetic digraph $\mathcal{D}$} of $\mathcal{N}$}, where the minimum is taken over all embeddings of $\mathcal{D}$} in $\mathcal{N}$} and their extensions, it is sufficient to consider only a single extension of $\mathcal{D}$}.
Phylogenetic digraphs. Let $D$ be a connected directed acyclic graph, and let $Y$ be a finite set. We say that $D$ is a leaf-labelled acyclic digraph on $Y$ if one of the following applies:
@@
    $|Y|=0$ and $D$ is the isolated vertex $\rho$,
  2. (ii)
    $|Y|=1$ and $D$ is the isolated vertex labelled with the element in $Y$, or
  3. (iii)
    $|Y|\geqslant 1$, $D$ has at most one vertex of in-degree zero and out-degree one, in which case this vertex is $\rho$, the leaves of $D$ have in-degree one and out-degree zero and are bijectively labelled with the elements in $Y$, and all other vertices of $D$ have in-degree zero and out-degree two, in-degree one and out-degree two, or in-degree two and out-degree one.
Similar to the leaf set of a phylogenetic network, we refer to $Y$ as the leaf set of $D$. Furthermore, ${\mathcal{L}}(D)$ denotes the leaf set of $D$. In contrast to a phylogenetic network, a leaf-labelled acyclic digraph $D$ may have more than one vertex with in-degree zero. Moreover, for a vertex $w$ in $D$ that has two parents $v$ and $v^{\prime}$, there does not necessarily exist a vertex $u$ such that there are edge-disjoint directed paths from $u$ to $v$ and from $u$ to $v^{\prime}$ in $D$.
Let $\mathcal{N}$} be a phylogenetic network on $X$ with root $\rho$, and let $D$ be a leaf-labelled acyclic digraph on $Y$ with $Y\subseteq X$. Recall that $D$ may or may not contain a vertex $\rho$ with in-degree zero and out-degree one. We say that $\mathcal{N}$} displays $D$ if there exists a subgraph of $\mathcal{N}$} that is isomorphic to $D$ up to suppressing vertices with in-degree one and out-degree one, in which case we call the subgraph an embedding $M$ of $D$ in $\mathcal{N}$} and view the edge set of $M$ as a subset of the edge set of $\mathcal{N}$}. More generally, for a collection ${\mathcal{D}}=\{D_{1},D_{2},\ldots,D_{k}\}$ of leaf-labelled acyclic digraphs, we say that $\mathcal{N}$} displays $\mathcal{D}$} if there exists an embedding $M_{i}$ of $D_{i}$ in $\mathcal{N}$} for each $i\in\{1,2,\ldots,k\}$ such that $M_{j}$ and $M_{j^{\prime}}$ are vertex disjoint for all distinct $j,j^{\prime}\in\{1,2,\ldots,k\}$, in which case we refer to ${\mathcal{M}}=\{M_{1},M_{2},\ldots,M_{k}\}$ as an embedding of $\mathcal{D}$} in $\mathcal{N}$}. Now let $\mathcal{M}$} be an embedding of $\mathcal{D}$} in $\mathcal{N}$}. Recalling that we allow tree vertices in $\mathcal{M}$} to have in-degree and out-degree one, we say that $\mathcal{M}$} is tree-child if each non-leaf vertex of $\mathcal{M}$} has a child that is a tree vertex or a leaf.
Let $\mathcal{N}$} be a phylogenetic network on $X$ with root $\rho$. Let ${\mathcal{D}}=\{D_{\rho},D_{1},D_{2},\ldots,D_{k}\}$ be a collection of leaf-labelled acyclic digraphs. Then $\mathcal{D}$} is called a phylogenetic digraph of $\mathcal{N}$} if the following three properties are satisfied:
  1. (i)
    the leaf sets ${\mathcal{L}}(D_{\rho}),{\mathcal{L}}(D_{1}),{\mathcal{L}}(D_{2}),\ldots,{\mathcal{L}}(D_{k})$ partition $X$ and $D_{\rho}$ is the only element in $\mathcal{D}$} that contains $\rho$,
@@
    there exists an embedding ${\mathcal{M}}=\{M_{\rho},M_{1},M_{2},\ldots,M_{k}\}$ of $\mathcal{D}$} in $\mathcal{N}$}.
@@
Extensions and root extensions. Let $\mathcal{N}$} be a phylogenetic network on $X$. Furthermore, let ${\mathcal{M}}=\{M_{\rho},M_{1},M_{2},\ldots,M_{k}\}$ be an embedding of a phylogenetic digraph ${\mathcal{D}}=\{D_{\rho},D_{1},D_{2},\ldots,D_{k}\}$ of $\mathcal{N}$}. We obtain an extension $\mathcal{R}$} of $\mathcal{D}$} in $\mathcal{N}$} from $\mathcal{M}$} by initially setting $\mathcal{R}$$\mathcal{M}$} and then repeatedly applying one of the following two operations until no further such operation is possible:
@@
By construction, observe that $\mathcal{R}$} contains exactly $k+1$ connected components and that there is a natural bijection between these components and the components in $\mathcal{D}$}. We therefore set ${\mathcal{R}}=\{R_{\rho},R_{1},R_{2},\ldots,R_{k}\}$ and call $R_{i}$ an extension of $D_{i}$ in $\mathcal{N}$} for each $i\in\{\rho,1,2,\ldots,k\}$. It follows from the construction of $\mathcal{R}$} that there is no vertex of out-degree two in $\mathcal{R}$} that is not also a vertex of out-degree two in $\mathcal{M}$}. Moreover, by construction, any underlying cycle in $\mathcal{R}$} is also an underlying cycle in $\mathcal{M}$}. Lastly, we define $\mathcal{R}$} to be tree-child precisely if $\mathcal{M}$} is tree-child, and refer to $\mathcal{M}$} as the embedding of $\mathcal{D}$} that underlies $\mathcal{R}$}.
@@
We next introduce a special type of extension. We call an extension $\mathcal{R}$} of $\mathcal{D}$} in $\mathcal{N}$} a root extension of $\mathcal{D}$} in $\mathcal{N}$} if it can be obtained from $\mathcal{M}$} by initially setting $\mathcal{R}$$\mathcal{M}$} and then repeatedly applying (E1) only until no further such operation is possible. Now let ${\mathcal{R}}=\{R_{\rho},R_{1},R_{2},\ldots,R_{k}\}$ be a root extension of $\mathcal{D}$} in $\mathcal{N}$}. Similar to the terminology for an extension, we call $R_{i}$ a root extension of $D_{i}$ in $\mathcal{N}$} for each $i\in\{\rho,1,2,\ldots,k\}$. Let $r$ be a vertex of in-degree zero and out-degree zero or two in $\mathcal{D}$}, and let $P$ be the unique maximal length directed path in $\mathcal{R}$} that starts at a vertex $u$ of in-degree zero and ends at $r$. We refer to $P$ as the root path of $r$. Note that $P$ may have no edge in which case $u=r$. If $u\neq r$, then $u$ has in-degree zero and out-degree one in $\mathcal{R}$}. Figure 2 illustrates the concepts of phylogenetic digraphs, extensions, and root extensions.
[…]
Figure 2: Two phylogenetic networks $\mathcal{N}$} and $\mathcal{N}$$^{\prime}$ and three phylogenetic digraphs ${\mathcal{D}}_{1}$, ${\mathcal{D}}_{2}$, and ${\mathcal{D}}_{3}$ for $\mathcal{N}$} and $\mathcal{N}$$^{\prime}$. For each $\in\{1,2,3\}, i{\mathcal{R}}_{i}$ is an extension of ${\mathcal{D}}_{i}$ in $\mathcal{N}$} and ${\mathcal{R}}_{i}^{\prime}$ is an extension of ${\mathcal{D}}_{i}$ in $\mathcal{N}$$^{\prime}$, where the edges and vertices of the extensions are indicated in red. Note that ${\mathcal{R}}_{3}$ is not a root extension of ${\mathcal{D}}_{3}$ in $\mathcal{N}$}.
@@
To ease reading throughout the remainder of the paper, we often consider a phylogenetic network $\mathcal{N}$}, a phylogenetic digraph $\mathcal{D}$} of $\mathcal{N}$}, and an extension $\mathcal{R}$} of $\mathcal{D}$} in $\mathcal{N}$}. In this case, we view the vertex and edge set of $\mathcal{R}$} (as well as the vertex and edge set of any embedding of $\mathcal{D}$} in $\mathcal{N}$}) as a subset of the vertex and edge set of $\mathcal{N}$}, respectively. Furthermore, for clarity, the in-degree (resp. out-degree) of a vertex $v$ in $\mathcal{R}$} refers to the number of edges in $\mathcal{R}$} that are directed into (resp. out of) $v$. Lastly, let ${\mathcal{D}}=\{D_{\rho},D_{1},D_{2},\ldots,D_{k}\}$ be a phylogenetic digraph of a phylogenetic network $\mathcal{N}$}. Then there exists an embedding ${\mathcal{M}}=\{M_{\rho},M_{1},M_{2},\ldots,M_{k}\}$ of $\mathcal{D}$} in $\mathcal{N}$} such that $M_{i}$ and $M_{j}$ are vertex disjoint for all distinct $i,j\in\{\rho,1,2,\ldots,k\}$ and an extension ${\mathcal{R}}=\{R_{\rho},R_{1},R_{2},\ldots,R_{k}\}$ of $\mathcal{D}$} in $\mathcal{N}$} such that $R_{i}$ and $R_{j}$ are vertex disjoint for all distinct $i,j\in\{\rho,1,2,\ldots,k\}$. It follows that each edge in $\mathcal{D}$} corresponds to a unique directed path in $\mathcal{M}$} (resp. $\mathcal{R}$}) whose non-terminal vertices all have in-degree one and out-degree one in $\mathcal{M}$} (resp. $\mathcal{R}$}), and each vertex in $\mathcal{D}$} corresponds to a unique vertex in $\mathcal{M}$} (resp. $\mathcal{R}$}). Reversely, each edge in $\mathcal{M}$} corresponds to a unique edge in $\mathcal{D}$}. We will freely use this correspondence throughout the paper.
4 Properties of extensions
@@
Refer to captionvvwwvvww(I)(II)$(E1)$ or (E2)vvwwvvww
Figure 3: Assuming that $v\notin{\mathcal{R}}$, the setup as described in the proof of Lemma 5(ii) for when (I) $w$ is a tree vertex or a leaf, and (II) $w$ is a reticulation in $\mathcal{N}$}. In both cases, an application of (E1) or (E2) can be used to extend $R_{i}$ by an additional edge so that the resulting extension contains $v$. Red solid lines indicate vertices and edges of $R_{i}$. Black (resp. red) dashed lines indicate vertices and edges of $\mathcal{N}$} (resp. $R_{i}$) that may or may not be vertices and edges of $\mathcal{N}$} (resp. $R_{i}$).
@@
Let ${\mathcal{D}}=\{D_{\rho},D_{1},D_{2},\ldots,D_{k}\}$, and let ${\mathcal{R}}=\{R_{\rho},R_{1},R_{2},\ldots,R_{k}\}$. To see that (i) holds, recall that $\mathcal{R}$} does not contain any vertex with out-degree zero that is not in $X\cup\{\rho\}$. To complete the proof, we establish (ii). This part of the proof is illustrated in Figure 3. By definition of $\mathcal{D}$}, the root and each leaf of $\mathcal{N}$} is contained in $\mathcal{R}$}. Towards a contradiction, we may therefore assume that there is a tree vertex or a reticulation in $\mathcal{N}$} that is not contained in $\mathcal{R}$}. Let $v$ be a vertex of $\mathcal{N}$} that is not in $\mathcal{R}$} such that every vertex that is distinct from $v$ and lies on a directed path from $v$ to a leaf in $\mathcal{N}$} is in $\mathcal{R}$}. Furthermore, let $w$ be a child of $v$. If $w$ is a tree vertex or leaf in $\mathcal{N}$}, then there exists a component $R_{i}$ with $i\in\{\rho,1,2,\ldots,k\}$ such that $w$ has in-degree zero in $R_{i}$, thereby contradicting that $R_{i}$ is an extension of $D_{i}$ as we can apply (E1). Otherwise, if $w$ is a reticulation in $\mathcal{N}$}, then it follows from (i) that there exists a component $R_{i}$ with $i\in\{\rho,1,2,\ldots,k\}$ such that $w$ has in-degree zero, or in-degree one and out-degree one in $R_{i}$, thereby again contradicting that $R_{i}$ is an extension of $D_{i}$ as we can apply either (E1) or (E2), respectively. ∎
Lemma 6.
@@
Let $u$ be the parent of $v$, and let $w$ and $w^{\prime}$ be the two children of $v$ in $\mathcal{N}$}. Furthermore, let ${\mathcal{R}}=\{R_{\rho},R_{1},R_{2},\ldots,R_{k}\}$. We first show that (i) holds. Since $\mathcal{M}$} does not contain any vertex of out-degree zero that is not in $X\cup\{\rho\}$, it follows that at least one of $(v,w)$ and $(v,w^{\prime})$ is in $\mathcal{M}$}. Without loss of generality, we may assume that $(v,w)\in{\mathcal{M}}$. The setup of the proof is shown in Figure 4(i). If $(v,w^{\prime})\in{\mathcal{M}}$, then the result clearly holds. On the other hand, if $(v,w^{\prime})\notin{\mathcal{M}}$, then it follows from the definition of an extension that $(v,w^{\prime})\notin{\mathcal{R}}$, thereby establishing (i). We now turn to (ii) which is illustrated in Figure 4(ii). It follows from Lemma 5(ii) that $v\in{\mathcal{R}}$. Since $\mathcal{R}$} does not contain any vertex with out-degree zero that is not in $X\cup\{\rho\}$, at least one of $(v,w)$ and $(v,w^{\prime})$ is contained in $\mathcal{R}$}. Moreover, again by the definition of an extension, at most one of $(v,w)$ and $(v,w^{\prime})$ is contained in $\mathcal{R}$}. The lemma now follows. ∎
[…]
Figure 4: Setup as described in the proof of parts (i) and (ii) of Lemma 6. The red solid line in (i) (resp. (ii)) indicates that $(v,w)$ is an edge of $\mathcal{M}$} (resp. $\mathcal{R}$}), and the red dashed line in (i) indicates that $(v,w^{\prime})$ may or may not be an edge of $\mathcal{M}$}.
@@
Let $\mathcal{M}$} and $\mathcal{M}$$^{\prime}$ be the embeddings of $\mathcal{D}$} in $\mathcal{N}$} that underlie $\mathcal{R}$} and $\mathcal{R}$$^{\prime}$, respectively. First assume that $\mathcal{M}$$\mathcal{M}$$^{\prime}$. Since $|E_{\mathcal{N}}-E_{\mathcal{M}}|=|E_{\mathcal{N}}-E_{{\mathcal{M}}^{\prime}}|$ and each iteration of (E1) or (E2) in the construction of $\mathcal{R}$} and $\mathcal{R}$$^{\prime}$ either adds a new edge $(v,w)$ such that $w$ is already in $\mathcal{R}$} and $v$ is not already in $\mathcal{R}$} or a new edge $(v^{\prime},w^{\prime})$ such that $w^{\prime}$ is already in $\mathcal{R}$$^{\prime}$ and $v^{\prime}$ is not already in $\mathcal{R}$$^{\prime}$, the result follows from Lemma 5(ii). Second assume that $\mathcal{M}$$\neq{\mathcal{M}}^{\prime}$. Consider an edge $e=(v,w)$ of $\mathcal{N}$}. If $v=\rho$, then, as $\rho$ has out-degree one in $\mathcal{N}$}, either $e\in{\mathcal{R}}$ and $e\in{\mathcal{R}}^{\prime}$, or $e\notin{\mathcal{R}}$ and $e\notin{\mathcal{R}}^{\prime}$. Furthermore, if $v$ is a reticulation in $\mathcal{N}$}, then it follows from Lemma 5(i) that $e\in{\mathcal{R}}$ and $e\in{\mathcal{R}}^{\prime}$. We next consider all edges in $\mathcal{N}$} that are directed out of a tree vertex. Let $v$ be a tree vertex of $\mathcal{N}$}, and let $e=(v,w)$ and $e^{\prime}=(v,w^{\prime})$ be the two edges directed out of $v$. We next consider four cases that will subsequently be used for a counting argument.